The perimeter of a rectangular sign is not to exceed 50 feet. The length is to be twice the width. What widths will meet these conditions?
The only formula I know is: P=2(l+w) I dont understand how this formula solves this problem.....please help me. Your time and efforts will be greatly appreciated. Thank You.

Answer by kingme18(98)

Since the length is to be twice the width, you can say l = 2w (length equals two times width). That means that l is the same as 2w...so in your formula, instead of l, you can say 2w: P = 2(2w + w)
Since perimeter is 50 ft, plug that in for P and solve:
50=2(2w+w)
50=2(3w)
50=6w
8.333...=w
w is 8.333... and l is twice that, or 16.666...
Since this is in terms of feet, we can write 8.333... ft as 8 ft, 4 inches (because .333... is 1/3, and 1/3 of a foot is 4 inches) and 16.666... ft as 16 ft, 8 inches.